OpenStudy (anonymous):

If AB=9X-1, and AC=3x+7, find the following x= AC=

OpenStudy (anonymous):

Is there more?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):



OpenStudy (anonymous):

If AC is equivalent to CB, then AB = 2(AC).

OpenStudy (anonymous):

Using AB = 2AC, sub in the expressions for AB and AC from the problem, then solve for x. At the end, use that x value to solve for AC

OpenStudy (anonymous):

I need help with the equation. Can you help?

OpenStudy (anonymous):

AB=9X-1 and AC=3x+7 The whole bottom of the triangle is AB, and it is twice as big as AC, or = 2AC So AB = 2 AC and using the top expressions, (9x - 1) = 2(3x + 7)

OpenStudy (anonymous):

Does that get you started?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

glad to help :)